# High Fidelity Text-Guided Music Generation and Editing via Single-Stage Flow Matching

### v1

@@ 6 Discussion
Under our current setup text-guided music editing prompts are not instructions. They describe what the edited sample should sound like given an original music sample and description, but the model is not designed to understand direct editing instructions like replace instrument A by instrument B. While we observed that MelodyFlow performs convincing editing tasks for several axes (genre or instrument swap, tempo modification, key transposition, inpainting/outpainting), more research work is required to accurately evaluate each of those axes. Music editing human listening tests are conducted for a fixed $T_{edit}$, but eventually it should depend on the sound designer’s preference on the creativity-consistency axis. Finally the reported objective metrics are mostly used as a proxy for subjective evaluations but they have their limitations. As an example we observe that optimizing FAD for MusicCaps is usually achieved by overfitting on our training dataset, which negatively correlates with perceived quality.
Broader Impact
Such a model has the potential to help creators expand their creativity by generating different styles of high quality music. The combination of efficiency and quality may eventually revolutionise the music sound design and composition work. This should unlock the ability for any user to listen to the music they like under a different style. However such a model can also give malicious users the ability to alter and republish any copyrighted content without the original creators content.
@@ A.1.1 Audio latent representation
Our compression model implementation is that of Copet et al. [2024] enhanced by band-wise discriminators and snake activations from [Kumar et al., 2024], perceptual weighting [Wright and Välimäki, 2019], VAE bottleneck and multi resolution STFT reconstruction loss from [Evans et al., 2024a]. We train a mono 32 kHz codec at 20 Hz frame rate and another one supporting stereo 48 kHz audio at 25 Hz. The bottleneck dimension is of 128. Both are trained on one-second random audio crops for 200K steps, with a constant learning rate of 0.0003, AdamW optimizer and loss balancer of [Défossez et al., 2022]. Stereo codecs are trained with sum and difference loss [Steinmetz et al., 2020]. The bottleneck layer statistics are tracked during training (dimension-wise) for normalization prior to FM model training.
@@ A.1.2 Flow matching model
MelodyFlow is a diffusion transformer that follows Esser et al. [2024] configurations where each head dimension is of 64 and the model has the same number of heads and layers (either 18 or 24). Model implementation is that of audiocraft [^3] https://github.com/facebookresearch/audiocraft but adapted for FM following [Vyas et al., 2023]: U-shaped skip connections are added along with linear projections applied after concatenation with each transformer block output Bao et al. [2023]. The model is conditioned via cross attention on a T5 representation [Raffel et al., 2020] computed from the text description of the audio, using 20% dropout rate during training in anticipation for the classifier free guidance applied at inference. Cross attention masking is used to properly adapt to the text conditioning sequence length of each sample within a batch and we use zero attention for the model to handle unconditional generation transparently. No prepossessing is applied on text data and we only rely on the descriptions (additional annotations tags such as musical key, tempo, type of instruments, etc. are discarded, although they also sometimes appear in the text description). The flow step is injected following [Hatamizadeh et al., 2023]. Minibatch coupling is computed with torch-linear-assignement [^4] https://github.com/ivan-chai/torch-linear-assignment. MelodyFlow-small (resp. MelodyFlow-medium) is trained on latent representation sequences of 32 kHz mono (resp. 48 kHz stereo) segments of 10 (resp. 30) seconds, encoded at 20 Hz frame rate (resp. 25 Hz). MelodyFlow-small (resp. MelodyFlow-medium) is trained for 240k (resp. 120k) steps with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay of 0.1 and gradient clipping at 0.2), a batch size of 576 and a cosine learning rate schedule with 4000 warmup steps. Additionally, we update an exponential moving average of the model weights ever 10 steps with a decay of 0.99. Each model is trained on 8 H100 96GB GPUs with bfloat16 mixed precision and FSDP [Zhao et al., 2023]. MelodyFlow-small requires 3 days and MelodyFlow-medium 6 days of training.
@@ A.1.3 LLM-assisted editing prompt generation
For editing prompts design we utilized the LLama-3 large language model to modify the original descriptions by targeting genre swapping. Edited descriptions were then manually verified to ensure their plausibility and coherence. As an example, given the original description This is a lush indie-folk song featuring soaring harmony interplay and haunting reverb-y harmonica, the resulting editing prompt is This is a lush Indian classical-inspired song featuring soaring harmony interplay and haunting reverb-y bansuri flute.
@@ A.2 Additional experiments
A.2.1 Classifier-free guidance
In the Figure 4(a) we report the in-domain evaluation FAD as a function of the classifier-free guidance factor in the text-to-music generation setting. Throughout the paper we use a classifier-free guidance factor of $5.0$.
A.2.2 Inference efficiency
In the Figure 4(b) we present the measured FAD as a function of inference steps for both the baseline and final version of MelodyFlow. Not only is MelodyFlow achieving better performance overall but it is able to outperform the baseline with 16 times fewer NFE.
(a) Classifier-free guidance.
@@ A.2.2 Inference efficiency
Figure 4: Text-to-music generation quality (FAD) as a function of the classifier-free guidance factor (Figure 4(a)) and inference steps (Figure 4(b)). The baseline of the Figure 4(b) is the FM model architecture of [Le et al., 2024] but retrained on our music latents. The combination of our flow matching design choices enable faster generation for a given efficiency budget or better overall quality.
A.2.3 Regularized inversion ablation
During latent inversion the FM model predicts the velocity $v_{\Theta}(\mathbf{z}_{t},t,c)=\mathbf{x}-\mathbf{\epsilon}$. The predictions are regularized using the Algorithm 1 using a weighted KL patch-wise divergence loss $\mathcal{L}_{patchKL}$. In the Figures 5(a) and 5(b) we ablate on the divergence loss weight $\lambda_{KL}$ for $T_{edit}=0.04$, using $S=25,K=4,w_{k}=k-1$. Since during latent inversion we know the original latent $\mathbf{x}_{orig}$, the model prediction can be rewritten as $\epsilon$-prediction by computing $\epsilon=v_{\Theta}(\mathbf{z}_{t},t,c)-\mathbf{x}_{orig}$. In that scenario we can apply the exact ReNoise inversion method of [Garibi et al., 2024], using 10 iterations of noise de-correlation (shown as $\epsilon$-prediction in the Figures). Whether the model prediction is expressed as noise or velocity, the Figures show that an optimum can be achieved around $\lambda_{KL}=0.15$ for velocity prediction and around $0.05$ for noise prediction. Overall the quality is better (lower FADedit in the Figure 5(a)) when directly regularizing the velocity prediction. In both cases we observe a higher CLAPedit in the Figure 5(b) when the original text description $c_{orig}$ conditions the inversion process, confirming better text-adherence. This happens at the expense of a higher FADedit compared with unconditional inversion.
(a) FADedit as a function of $\lambda_{KL}$.
(b) CLAPedit as a function of $\lambda_{KL}$.
Figure 5: Effect of the loss regularization weight $\lambda_{KL}$ on the quality (Figure 5(a)) and text-adherence (Figure 5(b)) of music editing. Noise and velocity prediction are compared, with or without the original text description $c_{orig}$.

### v2

@@ 6 Discussion
The proposed model specifically focuses on text-guided audio editing with the quality/efficiency trade off in mind, hence we do to not aim nor claim to outperform previous state of the art text-to-music generation models. Under our current setup text-guided music editing prompts are not instructions. They describe what the edited sample should sound like given an original music sample and description, but the model is not designed to understand direct editing instructions like replace instrument A by instrument B. While we observed that MelodyFlow performs convincing editing tasks for several axes (genre or instrument swap, tempo modification, key transposition, inpainting/outpainting), more research work is required to accurately evaluate each of those axes. Music editing human listening tests are conducted for a fixed $T_{edit}$, but eventually it should depend on the sound designer’s preference on the creativity-consistency axis. Finally the reported objective metrics are mostly used as a proxy for subjective evaluations but they have their limitations. As an example we observe that optimizing FAD for MusicCaps is usually achieved by overfitting on our training dataset, which negatively correlates with perceived quality. Overall subjective evaluations remain the best source of truth until a model that mimics human ratings is developed.
@@ A.1.1 Audio latent representation
Our compression model implementation is that of Copet et al. (2024) [^3] https://github.com/facebookresearch/audiocraft/blob/main/audiocraft/models/encodec.py enhanced by band-wise discriminators and snake activations from Kumar et al. (2024), perceptual weighting (Wright & Välimäki, 2019), VAE bottleneck and multi resolution STFT reconstruction loss from Evans et al. (2024a). We train a mono 32 kHz codec at 20 Hz frame rate and another one supporting stereo 48 kHz audio at 25 Hz. The bottleneck dimension is of 128. Both are trained on one-second random audio crops for 200K steps, with a constant learning rate of 0.0003, AdamW optimizer and loss balancer of (Défossez et al., 2022). Stereo codecs are trained with sum and difference loss (Steinmetz et al., 2020). The bottleneck layer statistics are tracked during training (dimension-wise) for normalization prior to FM model training.
@@ A.1.2 Flow matching model
MelodyFlow’s DiT follows Esser et al. (2024) configurations where each head dimension is of 64 and the model has the same number of heads and layers (either 18 or 24). Model implementation is that of audiocraft [^4] https://github.com/facebookresearch/audiocraft/blob/main/audiocraft/modules/transformer.py but adapted for FM following Vyas et al. (2023): U-shaped skip connections are added along with linear projections applied after concatenation with each transformer block output Bao et al. (2023). The model is conditioned via cross attention on a T5 representation (Raffel et al., 2020) computed from the text description of the audio, using 20% dropout rate during training in anticipation for the classifier free guidance applied at inference. Cross attention masking is used to properly adapt to the text conditioning sequence length of each sample within a batch and we use zero attention for the model to handle unconditional generation transparently. No prepossessing is applied on the text data and we only rely on the descriptions (additional annotations tags such as musical key, tempo, type of instruments, etc. are discarded, although they also sometimes appear in the text description). The flow timestep is injected following Hatamizadeh et al. (2023). Minibatch coupling is computed with torch-linear-assignement [^5] https://github.com/ivan-chai/torch-linear-assignment. MelodyFlow-small (resp. MelodyFlow-medium) is trained on latent representation sequences of 32 kHz mono (resp. 48 kHz stereo) segments of 10 (resp. 30) seconds, encoded at 20 Hz frame rate (resp. 25 Hz). MelodyFlow-small (resp. MelodyFlow-medium) is trained for 240k (resp. 120k) steps with AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay of 0.1 and gradient clipping at 0.2), a batch size of 576 and a cosine learning rate schedule with 4000 warmup steps. Additionally, we update an exponential moving average of the model weights ever 10 steps with a decay of 0.99. Each model is trained on 8 H100 96GB GPUs with bfloat16 mixed precision and FSDP (Zhao et al., 2023). MelodyFlow-small requires 3 days and MelodyFlow-medium 6 days of training.
@@ A.1.3 LLM-assisted editing prompt generation
For editing prompts design we prompted the LLama-3 large language model Dubey et al. (2024) to modify the original descriptions by targeting genre swapping. Edited descriptions were then manually verified to ensure their plausibility and coherence. As an example, given the original description This is a lush indie-folk song featuring soaring harmony interplay and haunting reverb-y harmonica, the resulting editing prompt is This is a lush Indian classical-inspired song featuring soaring harmony interplay and haunting reverb-y bansuri flute.
@@ A.2 Additional experiments
A.2.1 Stereo codec
The Table 5 reports the impact of scaling from mono to stereo with the same MelodyFlow-medium model size (1B parameters) trained on 30 second segments. Two codecs are trained on 48 kHz audio, using the same 25 Hz latent frame rate: one mono and one stereo. The in-domain FAD is reported for 10s and 30s generated segments. Moving from mono to stereo marginally affects the generative model performance.
A.2.2 Length generalization
One drawback of training the FM model of a fixed segment duration is that the inference can only be run for the same duration, otherwise the quality will degrade (this can be seen in the FAD10s column of the Table 5, when comparing the first two rows). This can be handled by using padded segments and specific conditioning (Evans et al., 2024a), but does not save any resource when targeting shorter segments. Another solution is to train on variable length segments but then the model does not generalize well for full length segments, and will better learn for the uttermost left positions of the sequence that appear more often. We propose to simulate training on variable length segments, while keeping the model learning for the full length scenario. This is done by applying a L-shaped attention mask during model. For each sequence of length L, we randomly select a segment boundary within the range $[0,L]$. Positions before the boundary can only attend to themselves in the DiT’s self-attention, while positions after it attend to the entire sequence.
Comparing the first two rows of the Table 5 indicate that the L-shaped mask helps supporting versatile duration with no penalty on full-length segments, unlocking faster inference for segments shorter than 30 seconds. This method does not generalize to segments longer than 30 seconds, which should be specifically handled with a sliding window/outpainting approach.
Table 5: Ablation on L-shaped attention mask and stereo for MelodyFlow-large. Each variant is trained on 30s audio segments encoded with a 25 Hz frame rate codec trained on 48 kHz audio.
[…]
2 0.40 12.48 0.59 0.65
1.48 0.65
1 0.39 13.34 0.49 0.59
(a) Generation CFG.
@@ A.2.2 Length generalization
Figure 6: Text-to-music generation quality (FAD) as a function of classifier-free guidance factor (Figure 6(a)) and inference steps (Figure 6(b)). The baseline of the Figure 6(b) is the FM model architecture of (Le et al., 2024) but retrained on our music latents. The combination of our flow matching design choices enable faster generation for a given efficiency budget or better overall quality.
A.2.3 Text-to-music generation efficiency
In the Figure 6(b) we report the text-to-music generation test FAD as a function of the number of DiT forward passes (NFEs) for both the baseline FM architecture (Le et al. (2024)) and final version of MelodyFlow. Not only does MelodyFlow achieve better performance, but with 16 times fewer NFEs (e.g. where the baseline required 256 NFEs to reach 0.53 FAD, MelodyFlow only requires 16 NFEs to score 0.50).
(a) Editing FAD vs. CFG.
(b) Editing CLAP vs. CFG.
(c) Editing LPAPS vs. CFG.
Figure 7: Music editing objective metrics as a function of the classifier free guidance, using the same CFG for both inversion and regeneration.
(a) Editing FAD vs. CFG.
(b) Editing CLAP vs. CFG.
(c) Editing LPAPS vs. CFG.
Figure 8: Music editing objective metrics as a function of the classifier free guidance, when using a CFG of 0 during latent inversion.
A.2.4 Classifier-free guidance
In the Figure 6(a) we report the in-domain test FAD as a function of the classifier-free guidance factor in the text-to-music generation setting. We use a classifier-free guidance factor of $4.0$ for the text-to-music generation inference.
In the Figures 7(a), 7(b), 7(c) we plot our objective metrics for text-guided music editing, as a function of the CFG. The performance is bad whatever the considered inversion method, showing that using the CFG during inversion is detrimental. Above a CFG of 0 ReNoise completely diverges (the LPAPS skyrockets), while MelodyFlow achieves the best robustness. This explains why we consider that ReNoise is not directly compatible with the FM formulation, even after converting to $\epsilon$-prediction, and that FM requires an adapted regularized latent inversion method. After keeping the CFG to zero during latent inversion to stabilize the process, the results are presented in the Figures 8(a), 8(b), 8(c). We end up using the same classifier-free guidance factor of $4.0$ for our editing experiments.